If f(x)=x^2+5 find f(a+h)-f(a)
Answer by stanbon(48558)

If f(x)=x^2+5 find f(a+h)-f(a)
--------------
f(a+h)= (a+h)^2+5
=a^2+2ah+h^2+5
f(a)=a^2+5
Then f(a+h)-f(a)= 2ah+h^2
Cheers,
Stan H.